Solve the following in the interval [0;2 π]

sinx > √3/2

cosx > √3/2

IcosxI ≤ 1/2


1
Expert's answer
2021-09-13T12:59:23-0400

(i)

the value of sin is +ve in first and second quadrant so the interval where sinx is positve is "0\space to\space \pi"

now we want sinx to be greater then "\frac {\sqrt3}{2}" ,now in first quadrant the graph of sin increases from "o\space to\space \frac{\pi}{2}"

and "sin\frac{\pi}{3}" is "\frac {\sqrt3}{2}" so in the interval "\frac{\pi}{3}\space to\space \frac{\pi}{2}" sin should be greater than "\frac {\sqrt3}{2}"

now in second quadrant the graph of sin decreases since the value of "sin\frac{\pi}{2}" is 1 and the value of "sin\frac{2\pi}{3}" is "\frac {\sqrt3}{2}" so the total interval in which sinx is greater than "\frac {\sqrt3}{2}\space is\space (\frac{\pi}{3}\space to\space \frac{2\pi}{3})"

(ii)

the value of cos is +ve in first and fourth quadrant so the interval where cosx is positve is "0\space to\space \frac{\pi}{2} \space and\space\frac{3\pi}{2}\space to\space 2\pi" since cos0 is 1 and cosx is decreasing graph in first quadrant and "cos\frac{\pi}{6}\space is \space \frac {\sqrt3}{2}"

so in first quadrant cosx greater than "\frac {\sqrt3}{2}" in interval 0 to "\frac{\pi}{6}"

now let's take fourth quadrant the value of cos increase in fourth quadrant since "cos\frac{11\pi}{6} \space is \space \frac {\sqrt3}{2}" and "cos2\pi" is 1, so the interval in fourth quadrant is "\frac{11\pi}{6}" to "2\pi"

so the total interval in which cosx is greater than "\frac {\sqrt3}{2}\space is\space (0\space to\space \frac{\pi}{6}) \cup (\frac{11\pi}{6}\space to \space 2\pi)"

(iii)

let's take first quadrant graph of cosx is decrease in first quadrant and the value of "cos\frac{\pi}{6}\space is \space \frac{1}{2}" since we need value less than equal to "\frac{1}{2}" so the required interval will be "[\frac{\pi}{3}" , "\frac{\pi}{2}]"

Now let's move forward to second quadrant the graph of cosx is still decrease in second quadrant but we have |cosx| the graph will be increasing the value of "|cos\frac{2\pi}{3}|=\frac{1}{2}" so the required interval is "[\frac{\pi}{2}\space ,\space \frac{2\pi}{3}]"

Now let's move forward to 3rd quadrant, graph of cosx increase is 3rd quadrant but we have |cosx| so the graph will be decreasing and the value of "|cos\frac{4\pi}{3}|=\frac{1}{2}" so the required interval is "[\frac{4\pi}{3}\space , \space \frac{3\pi}{2}]"

Now let's move forward to 4th quadrant in this the value of cosx is positive so |cosx| will not effect so the value of "|cos\frac{5\pi}{3}|=\frac{1}{2}" so the required interval is "[\frac{3\pi}{2}\space ,\space \frac{5\pi}{3}]"

so the total interval in which cosx is less than equal to "\frac{1}{2} \space [\frac{\pi}{3} \space ,\space \frac{2\pi}{3}] \cup [\frac{4\pi}{3}\space , \space \frac{5\pi}{3} ]"